OpenStudy (anonymous):

How much current does a 10 kΩ resistor draw from a 12.0 V battery? A. 1.2 × 10^-3 A B. 1.2 × 10^0 A C. 1.2 × 10^2 A D. 1.2 × 10^3 A

OpenStudy (michele_laino):

here is your computation: \[I = \frac{{\Delta V}}{R} = \frac{{12}}{{{{10}^4}}}\] since 10 kOhm= 10^4 ohms

OpenStudy (michele_laino):

\[\Large I = \frac{{\Delta V}}{R} = \frac{{12}}{{{{10}^4}}}\]

OpenStudy (anonymous):

ok! so we get 0.0012? would that mean choice A is our solution?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

ooh yay! thanks!:)

OpenStudy (michele_laino):

:)